OpenStudy (anonymous):

You must: Label and display your new polynomial identity Prove that it is true through an algebraic proof, identifying each step Demonstrate that your polynomial identity works on numerical relationships WARNING! No identities used in the lesson may be submitted. Create your own. See what happens when different binomials or trinomials are combined. Below is a list of some sample factors you may use to help develop your own identity.

OpenStudy (anonymous):

sample factors I can use.... (x – y) (x + y) (y + x) (y – x) (x + a) (y + b) (x^2 + 2xy + y^2) (x^2 – 2xy + y^2) (ax + b) (cy + d)

OpenStudy (anonymous):

from the warning not to use any in the lesson, I'd say just stay with the given samples.

ganeshie8 (ganeshie8):

just cookup some identity

ganeshie8 (ganeshie8):

maybe try to derive a formula for directly calculating :- (ax + b)(cx + d)

OpenStudy (anonymous):

I'm back, and I'm kind of confused on that lol

OpenStudy (anonymous):

An identity is like breaking down a equation so that you end up with a number being equal to the same number right?

ganeshie8 (ganeshie8):

did u get wat the question is asking us to do ? :)

OpenStudy (anonymous):

I need to create a identity and prove it through algebra I guss lol

ganeshie8 (ganeshie8):

An identity can be a formula, that u can use to save time.

OpenStudy (anonymous):

so I need to make up a formula lol?

ganeshie8 (ganeshie8):

for ex, below is an identity :- (a + b)(a - b) = a^2 - b^2

OpenStudy (anonymous):

lol idk what to make

ganeshie8 (ganeshie8):

read my second reply

ganeshie8 (ganeshie8):

make a formula for that

OpenStudy (anonymous):

(ax + b)(cx + d) how about... acx + bd?

ganeshie8 (ganeshie8):

(ax + b)(cx + d) FOIL it first, then try to simplify

OpenStudy (anonymous):

umm (ax + b)(cx + d) acx^2 + adx + bcx +bd

OpenStudy (anonymous):

brb real quick, lunch...

ganeshie8 (ganeshie8):

yes, we're done :- just take out x common

ganeshie8 (ganeshie8):

(ax + b)(cx + d) FOILing it :- acx^2 + adx + bcx +bd taking x common in middle two terms :- acx^2 + (ad+bc)x + bd

ganeshie8 (ganeshie8):

so ur new polynomial identity is :- (ax+b)(cx+d) = acx^2 + (ad+bc)x + bd

ganeshie8 (ganeshie8):

Enjoy ur lunch...

OpenStudy (anonymous):

ok I'm back

OpenStudy (anonymous):

Hey can you help me use this with a numerical equation now?

OpenStudy (anonymous):

@ganeshie8